A chemist wants to make a 50ml of a 16% acid solution by mixing a 13% acid solution and an 18% acid solution. how many milliliters of each solution should the chemist use.
.13(50-x)+.18x= 8
6.5-13x+.18x=8
6.5+5x=8
-6.5
5x=1.5
x=3
What am I doing wrong with these charts? i am guessing I am putting everything in the wrong place???

Answer by Alan3354(69443)

A chemist wants to make a 50ml of a 16% acid solution by mixing a 13% acid solution and an 18% acid solution. how many milliliters of each solution should the chemist use.
.13(50-x)+.18x= 8
6.5-13x+.18x=8
6.5+5x=8
-6.5
5x=1.5
x=3
What am I doing wrong with these charts? i am guessing I am putting everything in the wrong place???
----------------
Think about what the numbers mean.
----------------
t = amount of 13%
e = amount of 18%
--------
t + e = 50 (total solution)
18e + 13t = 50*16 (total acid)
---
Now it's 2 eqns in 2 variables